. From a box of a dozen rocket-propelled grenade (RPG), 4 are selected at random and fired. If the box contains 3 defective RPG’s, what is the probability that (a) all 4 will fire? (b) at most 2 will not fire?m


1
Expert's answer
2021-05-07T09:47:17-0400

Solution:

Given, "n=4,p=\frac9{12}=\frac34,q=\frac14"

Here, p represents the success of RPG will fire.

"X\sim Bin(n,p)"

(a): "P(X=4)=^4C_4(\frac3{4})^4(\frac1{4})^0=0.3164"

(b): "P(X\le2)=1-P(X>2)=1-[P(X=3)+P(X=4]"

"=^4C_3(\frac3{4})^3(\frac1{4})^1+^4C_4(\frac3{4})^4(\frac1{4})^0"

"=0.7382"

Now, the probability that at most two will not fire"=1-0.7382=0.2618"